# RAG vs. GraphRAG: A Systematic Evaluation and Key Insights

### v1

@@ Abstract
Haoyu Han1, Yu Wang2, Harry Shomer1, Yongjia Lei, 2, Kai Guo 1, Zhigang Hua3, Bo Long 3, Hui Liu1, Jiliang Tang1 1Michigan State University, 2University of Oregon, 3Meta {hanhaoy1, shomerha, guokai1, liuhui7, tangjili}@msu.edu {yuwang, yongjia}@uoregon.edu, {zhua, bolong}@meta.com
@@ 1 Introduction
Graphs, as a fundamental data structure, encode rich relational information and have been extensively utilized across real-world domains, including knowledge representation, social network analysis, and biomedical research Wu et al. (2020); Ma and Tang (2021); Wu et al. (2023). Motivated by this, GraphRAG has recently gained attention for retrieving graph-structured data, such as knowledge graphs (KGs) and molecular graphs (Han et al., 2024; Peng et al., 2024). Beyond leveraging existing graphs, GraphRAG has also demonstrated its effectiveness for text-based tasks after structuring implicit knowledge from text into graph representations, benefiting applications such as global summarization Edge et al. (2024); Zhang et al. (2024), planning Lin et al. (2024) and reasoning Han et al. (2025).
While previous studies have demonstrated the potential of GraphRAG for text-based tasks by converting sequential text into graphs, most of them primarily focus on specific tasks and well-designed datasets. Consequently, the applicability of GraphRAG to broader, real-world text-based tasks remains unclear, particularly when compared to RAG, which has seen widespread adoption across diverse applications. This raises a critical question: What are the advantages and disadvantages of applying GraphRAG to general text-based tasks compared to RAG?
@@
To bridge this gap, we systematically evaluate the performance of RAG and GraphRAG on general text-based tasks using widely adopted datasets, including Question Answering and Query-based Summarization. Specifically, we assess two representative GraphRAG methods: (1) Knowledge Graph-based GraphRAG Liu (2022), which extracts a Knowledge Graph (KG) from text and performs retrieval solely based on the KG and (2) Community-based GraphRAG (Edge et al., 2024), which retrieves information not only from the constructed KG but also from hierarchical communities within the graph. For the Question Answering task, we conduct experiments on both single-hop and multi-hop QA under single-document and multi-document scenarios. Similarly, for the Query-based Summarization task, we evaluate both single-document and multi-document summarization to comprehensively assess the effectiveness of RAG and GraphRAG.
@@
Based on our comprehensive evaluation, we conduct an in-depth analysis of the strengths and weaknesses of RAG and GraphRAG across different tasks. Our findings reveal that RAG and GraphRAG are complementary, each excelling in different aspects. For the Question Answering task, we observe that RAG performs better on single-hop questions and those requiring detailed information, while GraphRAG is more effective for multi-hop questions. In the Query-based Summarization task, RAG captures fine-grained details, whereas GraphRAG generates more diverse and multi-faceted summaries. Building on these insights, we investigate two strategies from different perspectives to integrate their unique strengths and enhance the overall performance. Our main contributions are as follows:
  • Systematical Evaluation : This is the very first work to systematically evaluate and compare RAG and GraphRAG on text-based tasks using widely adopted datasets and evaluations.
@@
  • Hybrid Retrieval Strategies: Based on our findings on the unique strengths of RAG and GraphRAG, we propose two strategies to improve overall performance: (1) Selection, where queries are dynamically assigned to either RAG or GraphRAG based on their characteristics, and (2) Integration, where both methods are integrated to leverage their complementary strengths.
  • Challenges and Future Directions: We discuss the limitations of current GraphRAG approaches and outline potential future research directions for broader applicability.
@@ 2.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) has been widely applied to enhance the performance of Large Language Models (LLMs) by retrieving relevant information from external sources, addressing the limitation of LLMs’ restricted context windows, improving factual accuracy, and mitigating hallucinations Fan et al. (2024); Gao et al. (2023). Most RAG systems primarily process text data by first splitting it into chunks Finardi et al. (2024). When a query is received, RAG retrieves relevant chunks either through lexical search Ram et al. (2023) or by computing semantic similarity Karpukhin et al. (2020), embeddings both the query and text chunks into a shared vector space. Advanced techniques, such as pre-retrieval processing Ma et al. (2023); Zheng et al. (2023a) and post-retrieval processing Dong et al. (2024); Xu et al. (2023), as well as fine-tuning strategies Li et al. (2023), have further enhanced RAG’s effectiveness across various domains, including QA) Yan et al. (2024), dialogue generation Izacard et al. (2023), and text summarization Jiang et al. (2023).
Several studies have evaluated the effectiveness of RAG systems across various tasks Yu et al. (2024); Chen et al. (2024); Es et al. (2023), such as multi-hop question answering Tang and Yang (2024), biomedical question answering Xiong et al. (2024), and text generation Liu et al. (2023). However, no existing study has simultaneously and systematically evaluated and compared RAG and GraphRAG on these general text-based tasks.
@@ 2.2 Graph Retrieval-Augmented Generation
Despite leveraging the existing graphs, recent studies have explored incorporating graph construction into GraphRAG to enhance text-based tasks. For example, Dong et al. (2024) construct document graphs using Abstract Meaning Representation (AMR) to improve document ranking. Edge et al. (2024) construct graphs from documents using LLMs, where nodes represent entities and edges capture relationships between them. Based on these graphs, they generate hierarchical communities and corresponding community summaries or reports. Their approach focuses on the global query summarization task, retrieving information from both the constructed graphs and their hierarchical communities. Additionally, Han et al. (2025) propose an iterative graph construction approach using LLMs to improve reasoning tasks.
@@ 3 Evaluation Methodology
In this section, we introduce the details of our evaluation framework. We primarily evaluate one representative RAG system and two representative GraphRAG systems, as illustrated in Figure 1.
@@ 3.1 RAG
We adopt a representative semantic similarity-based retrieval approach as our RAG method Karpukhin et al. (2020). Specifically, we first split the text into chunks, each containing approximately 256 tokens. For indexing, we use OpenAI’s text-embedding-ada-002 model, which has demonstrated effectiveness across various tasks Nussbaum et al. (2024). For each query, we retrieve chunks with Top-10 similarity scores. To generate responses, we employ two open-source models of different sizes: Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct Dubey et al. (2024).
For single-document tasks, we generate a separate RAG system for each document, ensuring that queries corresponding to a specific document are processed within its respective indexed chunk pool. For multi-document tasks, we use a shared RAG system by indexing all documents together.
@@ 3.2 GraphRAG
We select two representative GraphRAG methods for a comprehensive evaluation, as shown in Figure 1, namely KG-based GraphRAG and Community-based GraphRAG.
In the KG-based GraphRAG (KG-GraphRAG) Liu (2022), a knowledge graph is first constructed from text chunks using LLMs through triplet extraction. When a query is received, its entities are extracted and matched to those in the constructed KG using LLMs. The retrieval process then traverses the graph from the matched entities and gathers triplets (head, relation, tail) from their multi-hop neighbors as the retrieved content. Additionally, for each triplet, we can retrieve the corresponding text associated with it. We define two variants of KG-GraphRAG: (1) KG-GraphRAG (Triplets), which retrieves only the triplets, and (2) KG-GraphRAG (Triplets+Text), which retrieves both the triplets and their associated source text. We implement the KG-GraphRAG methods using LlamaIndex Liu (2022)  [^1] https://www.llamaindex.ai/.
For the Community-based GraphRAG Edge et al. (2024), in addition to generating KGs using LLMs, hierarchical communities are constructed using graph community detection algorithms, as shown in Figure 1. Each community is associated with a corresponding text summary or report, where lower-level communities contain detailed information from the original text. The higher-level communities further provide summaries of the lower-level communities. Due to the hierarchical community structure, there are two primary retrieval methods for retrieving relevant information given a query: Local Search and Global Search. In Local Search, entities, relations, their descriptions, and lower-level community reports are retrieved based on entity matching between the query’s extracted entities and the constructed graph. We refer to this method as Community-GraphRAG (Local). In Global Search, only high-level community summaries are retrieved based on semantic similarity to the query. We refer to this method as Community-GraphRAG (Global). The Community-GraphRAG methods are implemented using Microsoft GraphRAG Edge et al. (2024) [^2] https://microsoft.github.io/graphrag.
To ensure a fair comparison, we adopt the same settings for both RAG and GraphRAG methods. This includes the chunking strategy, embedding model, and LLMs. We select two representative RAG tasks, i.e., Question Answering and Query-based Summarization, to evaluate RAG and GraphRAG simultaneously.
@@ 4 Question Answering
QA is one of the most widely used tasks for evaluating the performance of RAG systems. QA tasks come in various forms, such as single-hop QA, multi-hop QA, and open-domain QA Wang (2022). To systematically assess the effectiveness of RAG and GraphRAG in these tasks, we evaluate them on widely used QA datasets and employ standard evaluation metrics.
Table 1: Performance comparison (%) on NQ and Hotpot datasets. The best results are highlighted in bold, and the second-best results are underlined.
@@
RAG 92.16 57.59 96.01 30.7 67.02 94.85 56.31 91.36 25.73 65.77
KG-GraphRAG (Triplets only) 55.76 22.55 98.67 18.7 41.24 76.96 32.36 94.35 19.55 50.98
KG-GraphRAG (Triplets+Text) 67.4 34.7 97.34 17.15 48.51 85.91 35.98 86.38 21.61 54.58
Community-GraphRAG (Local) 86.89 60.63 80.07 50.6 69.01 92.03 60.16 88.70 49.06 71.17
Community-GraphRAG (Global) 89.34 64.02 19.27 53.34 64.4 89.09 66.00 13.95 59.18 65.69
@@
Community-GraphRAG (Local) Community-GraphRAG (Global)
@@ 4.1 Datasets and Evaluation Metrics
To comprehensively evaluate the performance of GraphRAG on general QA tasks, we select four widely used datasets that cover different perspectives. For the single-hop QA task, we select the Natural Questions (NQ) dataset Kwiatkowski et al. (2019). For the multi-hop QA task, we select HotPotQA Yang et al. (2018) and MultiHop-RAG Tang and Yang (2024) datasets. The MultiHop-RAG dataset categorizes queries into four types: Inference, Comparison, Temporal, and Null queries. To further analyze the performance of RAG and GraphRAG at a finer granularity, we also include NovelQA Tang and Yang (2024), which contains 21 different types of queries. For more details, please refer to Appendix A.1.1. We use Precision (P), Recall (R), and F1-score as evaluation metrics for the NQ and HotPotQA datasets, while accuracy is used for the MultiHop-RAG and NovelQA datasets following their original papers.
@@ 4.2 QA Main Results
The performance comparison for the NQ and HotPotQA datasets is presented in Table 1, while that of MultiHop-RAG is shown in Table 2. Due to space constraints, partial results of NovelQA with the Llama 3.1-8B model are shown in Table 3, with the full results available in Appendix A.2. Based on these results, we make the following observations:
@@
    Community-GraphRAG (Global) often struggles on QA tasks. This is due to the global search retrieves only high-level communities, leading to a loss of detailed information. This is particularly evident from its lower performance on detail-oriented queries in the NovelQA dataset. Additionally, Community-GraphRAG (Global) tends to hallucinate in QA tasks, as shown by its poor performance on Null queries in the MultiHop-RAG dataset, which should ideally be answered as ‘insufficient information.’ However, this summarization approach may be beneficial for queries that require comparing different topics or understanding their temporal ordering, such as Comparison and Temporal queries in the MultiHop-RAG dataset, as shown in Table 2.
@@
    KG-based GraphRAG also generally underperform on QA tasks. This is because it retrieves information solely from the constructed knowledge graph, which contains only entities and their relations. However, the extracted entities and relations may be incomplete, leading to gaps in the retrieved information. To verify this, we calculated the ratio of answer entities present in the constructed KG. We found that only around 65.8% of answer entities exist in the constructed KG for the Hotpot dataset and 65.5% for the NQ dataset. These findings highlight a key limitation in KG-based retrieval and suggest the need for improved KG construction methods to enhance graph completeness for QA.
@@ 4.3 Comparative QA Analysis
The confusion matrices representing these four groups using the Llama 3.1-8B model are shown in Figure 2. Notably, the proportions of queries correctly answered exclusively by GraphRAG and RAG are significant. For example, 13.6% of queries are GraphRAG-only, while 11.6% are RAG-only on MultiHop-RAG dataset. This phenomenon highlights the complementary properties of RAG and GraphRAG, and each method has its own strengths and weaknesses. Therefore, leveraging their unique advantages has the potential to improve overall performance.

### v2

@@ Abstract
Haoyu Han1, Li Ma1 Yu Wang2, Harry Shomer1, Yongjia Lei, 2, Kai Guo 1, Zhigang Hua3, Bo Long 3, Hui Liu1, Charu C. Aggarwal4, Jiliang Tang1 1Michigan State University, 2University of Oregon, 3Meta, 4IBM T.J. Watson Research Center, {hanhaoy1, mali13, shomerha, guokai1, liuhui7, tangjili}@msu.edu {yuwang, yongjia}@uoregon.edu, {zhua, bolong}@meta.com, charu@us.ibm.com
@@ 1 Introduction
Graphs encode rich relational information and have been extensively utilized across real-world domains, including knowledge representation, social network analysis, and biomedical research Wu et al. (2020); Ma and Tang (2021); Wu et al. (2023). GraphRAG has recently gained attention for retrieving graph-structured data, such as knowledge graphs (KGs) and molecular graphs (Han et al., 2024; Peng et al., 2024). Beyond leveraging existing graphs, GraphRAG has also demonstrated its effectiveness for text-based tasks after structuring implicit knowledge from text into graph representations, benefiting applications such as global summarization Edge et al. (2024), planning Lin et al. (2024) and reasoning Han et al. (2025).
While recent studies have shown that constructing graphs from text can improve reasoning, planning, and summarization, most existing GraphRAG systems have been designed for narrow, task-specific settings with bespoke datasets and heuristics. As a result, the community still lacks a unified understanding of when and why explicit graph structures help or fail to help retrieval-augmented generation. This gap in understanding limits the principled use of GraphRAG beyond anecdotal success cases.
@@
To bridge this gap, we systematically evaluate the performance of RAG and GraphRAG on general text-based tasks using widely adopted datasets, including Question Answering and Query-based Summarization. Specifically, we assess three representative GraphRAG methods: (1) Knowledge Graph-based GraphRAG Liu (2022), which extracts a KG from text and performs retrieval solely based on the KG, (2) Community-based GraphRAG (Edge et al., 2024), which retrieves information not only from the constructed KG but also from hierarchical communities within the graph, and (3) Text-based GraphRAG (Jimenez Gutierrez et al., 2024), which retrieve original text chunks with the help of graph.
@@
Based on our comprehensive evaluation, we conduct an in-depth analysis of the strengths and weaknesses of RAG and GraphRAG across different tasks. Our findings reveal that RAG and GraphRAG are complementary, each excelling in different aspects. For the Question Answering task, we observe that RAG performs better on single-hop questions and those requiring fine-grained details, whereas GraphRAG is more effective for multi-hop and reasoning-intensive questions. In the Query-based Summarization task, RAG captures fine-grained details, whereas GraphRAG generates more diverse and multi-faceted summaries. Building on these insights, we investigate two strategies from different perspectives to integrate their unique strengths and enhance the overall performance. Our main contributions are as follows:
  • Systematical Evaluation: We present the first controlled, systematic evaluation of RAG and GraphRAG across both question answering and summarization, using identical LLMs, embeddings, and retrieval configurations. This unified setup enables a fair and replicable comparison across diverse text-based tasks.
@@
  • Hybrid Retrieval Strategies: We introduce two practical approaches, Selection and Integration, that adaptively combine RAG and GraphRAG to leverage their complementary strengths.
  • Challenges and Future Directions: Our analyses expose concrete bottlenecks, such as limited graph coverage and strong position bias in LLM-as-a-Judge evaluations, and outline paths toward more complete, efficient, and unbiased GraphRAG systems.
@@ 2.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) has been widely applied to enhance the performance of Large Language Models (LLMs) by retrieving relevant information from external sources, addressing the limitation of LLMs’ restricted context windows, improving factual accuracy, and mitigating hallucinations Fan et al. (2024); Gao et al. (2023). Most RAG systems primarily process text data by first splitting it into chunks Finardi et al. (2024). When a query is received, RAG retrieves relevant chunks either through lexical search Ram et al. (2023) or by computing semantic similarity Karpukhin et al. (2020), embeddings both the query and text chunks into a shared vector space. Advanced techniques, such as pre-retrieval processing Ma et al. (2023); Zheng et al. (2023a) and post-retrieval processing Dong et al. (2024); Xu et al. (2023), as well as fine-tuning strategies Li et al. (2023), have further enhanced RAG’s effectiveness across various domains, including QA) Yan et al. (2024), dialogue generation Izacard et al. (2023), and text summarization Jiang et al. (2023). Several studies have evaluated the effectiveness of RAG systems across various tasks Yu et al. (2024); Chen et al. (2024); Es et al. (2023), such as multi-hop question answering Tang and Yang (2024), biomedical question answering Xiong et al. (2024), and text generation Liu et al. (2023). However, no existing study has simultaneously and systematically evaluated and compared RAG and GraphRAG on these general text-based tasks.
@@ 2.2 Graph Retrieval-Augmented Generation
Despite leveraging the existing graphs, recent studies have explored incorporating graph construction into GraphRAG to enhance text-based tasks. For example, Dong et al. (2024) construct document graphs to improve document ranking. Edge et al. (2024) construct graphs from documents using LLMs, where nodes represent entities and edges capture relationships between them. Based on these graphs, they generate hierarchical communities and corresponding community summaries or reports. Their approach focuses on the global query summarization task, retrieving information from both the constructed graphs and their hierarchical communities. Additionally, Han et al. (2025) propose an iterative graph construction approach using LLMs to improve reasoning tasks.
@@ 3 Evaluation Methodology
In this section, we present the details of our evaluation framework. To ensure a fair comparison, we adopt the vanilla configurations for both RAG and GraphRAG under identical experimental settings.
@@ 3.1 RAG
We adopt a representative semantic similarity-based retrieval approach as our RAG method Karpukhin et al. (2020) as shown in Figure 1. Specifically, we first split the text into chunks, each containing approximately 256 tokens. For indexing, we use OpenAI’s text-embedding-ada-002 model, which has demonstrated effectiveness across various tasks Nussbaum et al. (2024). For each query, we retrieve chunks with Top-10 similarity scores. To generate responses, we employ two open-source models of different sizes: Llama-3.1-8B-Instruct and Llama-3.1-70B-Instruct Dubey et al. (2024).
@@ 3.2 GraphRAG
There are various designs of GraphRAG, with retrieval granularity typically falling into three main categories: (1) KG-based GraphRAG, which retrieves text-augmented triplets or raw triplets; (2) Community-based GraphRAG, which retrieves either local community subgraphs with triplets or global summaries; and (3) Text-based GraphRAG, which directly retrieves original text chunks using the graph structure. To ensure a comprehensive evaluation, we select representative methods from each category, as detailed below.
In the KG-based GraphRAG (KG-GraphRAG) Liu (2022), a knowledge graph is first constructed from text chunks using LLMs through triplet extraction. When a query is received, its entities are extracted and matched to those in the constructed KG using LLMs. The retrieval process then traverses the graph from the matched entities and gathers triplets (head, relation, tail) from their multi-hop neighbors as the retrieved content. Additionally, for each triplet, we can retrieve the corresponding text associated with it. We define two variants of KG-GraphRAG: (1) KG-GraphRAG (Triplets), which retrieves only the triplets, and (2) KG-GraphRAG (Triplets+Text), which retrieves both the triplets and their associated source text.
For the Community-based GraphRAG Edge et al. (2024), in addition to generating KGs using LLMs, hierarchical communities are constructed using graph community detection algorithms, as shown in Figure 1. Each community is associated with a corresponding text summary or report, where lower-level communities contain detailed information from the original text. The higher-level communities further provide summaries of the lower-level communities. Due to the hierarchical community structure, there are two primary retrieval methods for retrieving relevant information given a query: Local Search and Global Search. In Local Search, entities, relations, their descriptions, and lower-level community reports are retrieved based on entity matching between the query’s extracted entities and the constructed graph. We refer to this method as Community-GraphRAG (Local). In Global Search, only high-level community summaries are retrieved based on semantic similarity to the query. We refer to this method as Community-GraphRAG (Global). The Community-GraphRAG methods are implemented using Edge et al. (2024). In this paper, we primarily use GPT-4o-mini to construct the graphs. The results with GPT-4o are also provided in Appendix A.11.
For Text-based GraphRAG, the original text chunks are typically treated as nodes in the graph. We select HippoRAG2 Gutiérrez et al. (2025) as a representative method. HippoRAG2 constructs a KG from the original text chunks, where each entity is also connected back to the corresponding text chunks. During retrieval, entities relevant to the query are first identified, similar to KG-based GraphRAG, and then the original text chunks connected to these entities are retrieved.
To ensure a fair comparison, we adopt the same settings for both RAG and GraphRAG methods, including the chunking strategy, embedding model, and generation LLMs. We select two representative RAG tasks, i.e., Question Answering and Query-based Summarization, to evaluate RAG and GraphRAG simultaneously.
@@ 4 Question Answering
QA is one of the most widely used tasks for evaluating the performance of RAG systems. QA tasks come in various forms, such as single-hop QA, multi-hop QA, and open-domain QA Wang (2022). To systematically assess the effectiveness of RAG and GraphRAG in these tasks, we evaluate them on widely used QA datasets and evaluation metrics.
Table 1: Performance comparison (%) on NQ and Hotpot datasets.
@@
RAG 92.16 57.59 96.01 30.70 67.02 94.85 56.31 91.36 25.73 65.77
KG-GraphRAG (Triplets only) 55.76 22.55 98.67 18.70 41.24 76.96 32.36 94.35 19.55 50.98
KG-GraphRAG (Triplets+Text) 67.40 34.70 97.34 17.15 48.51 85.91 35.98 86.38 21.61 54.58
Community-GraphRAG (Local) 86.89 60.63 80.07 50.60 69.01 92.03 60.16 88.70 49.06 71.17
Community-GraphRAG (Global) 89.34 64.02 19.27 53.34 64.40 89.09 66.00 13.95 59.18 65.69
HippoRAG2 91.54 58.41 85.71 49.91 70.27 93.01 58.76 90.03 43.40 69.87
@@
Community-GraphRAG (Local) HippoRAG2
@@ 4.1 Datasets and Evaluation Metrics
To comprehensively evaluate the performance of GraphRAG on general QA tasks, we select four widely used datasets that cover different perspectives. For the single-hop QA task, we select the Natural Questions (NQ) dataset Kwiatkowski et al. (2019). For the multi-hop QA task, we select HotPotQA Yang et al. (2018) and MultiHop-RAG Tang and Yang (2024) datasets. The MultiHop-RAG dataset categorizes queries into four types: Inference, Comparison, Temporal, and Null queries. To further analyze the performance of RAG and GraphRAG at a finer granularity, we also include NovelQA Wang et al. (2024a), which contains 21 different types of queries. For more details, please refer to Appendix A.1.1. We use Precision (P), Recall (R), and F1-score as evaluation metrics for the NQ and HotPotQA datasets, while accuracy is used for the MultiHop-RAG and NovelQA datasets following their original papers.
@@ 4.2 QA Main Results
The performance comparison for the NQ and HotPotQA datasets is presented in Table 1, while that of MultiHop-RAG is shown in Table 2. The overall and average performance are reported as weighted averages. Due to space constraints, partial results of NovelQA with the Llama 3.1-8B model are shown in Table 3, with the full results available in Appendix A.2. Based on these results, we make the following observations:
@@
    Community-GraphRAG (Global) often struggles on QA tasks. This is due to the global search retrieves only high-level communities, leading to a loss of detailed information. This is particularly evident from its lower performance on detail-oriented queries in the NovelQA dataset. Additionally, Community-GraphRAG (Global) tends to hallucinate in QA tasks, as shown by its poor performance on Null queries in the MultiHop-RAG dataset, which should ideally be answered as ‘insufficient information.’ However, this summarization approach may be beneficial for queries that require comparing different topics or understanding their temporal ordering, such as Comparison and Temporal queries in the MultiHop-RAG dataset (Table 2).
@@
    KG-based GraphRAG also generally underperform on QA tasks. This is because it retrieves information solely from the constructed knowledge graph, which contains only entities and their relations. However, the extracted entities and relations may be incomplete, leading to gaps in the retrieved information. To verify this, we calculated the retrieval accuracy in Appendix A.3. We found that only around 65.8% of answer entities exist in the constructed KG for the Hotpot dataset and 65.5% for the NQ dataset. These findings highlight a key limitation in KG-based retrieval and suggest the need for improved KG construction methods to enhance graph completeness for QA.
We further provide case studies in Appendix A.4. Additionally, we compare the performance of RAG and GraphRAG under multi-step retrieval settings. The results, presented in the Appendix A.5, show consistent trends, further demonstrating the robustness of our conclusions.
@@ 4.3 Comparative QA Analysis
The confusion matrices representing these four groups using the Llama 3.1-8B model are shown in Figure 2. Notably, the proportions of queries correctly answered exclusively by GraphRAG and RAG are significant. For example, 13.6% of queries are GraphRAG-only, while 11.6% are RAG-only on MultiHop-RAG dataset. This phenomenon highlights the complementary properties of RAG and GraphRAG. Therefore, leveraging their unique advantages has the potential to improve overall performance.